OpenStudy (anonymous):

What volume is neeeded to have 0.1 m NaOH in a 0.05 M solution. Show work please. Formula is M=mol/L

OpenStudy (callisto):

0.1 m NaOH = 0.1mole NaOH?

OpenStudy (callisto):

if yes 0.1mol / vol required = 0.05M vol required = 0.1/0.5 = 0.2L

OpenStudy (anonymous):

Callisto, I have a question. If you have 0.1 mols of compound per 1.2 L of liquid, wouldn't that be 0.0833 mols/L?

OpenStudy (callisto):

yes

OpenStudy (anonymous):

I guess I'm reading the question wrong :-(

OpenStudy (anonymous):

I thought it meant changing the concentration from 0.1 M to 0.05M

OpenStudy (callisto):

ouch.. revised: if yes 0.1mol / vol required = 0.05M vol required = 0.1/0.05 = 2L

OpenStudy (anonymous):

Oh okay. I guess it's a good thing that even though I don't know what I'm doing, that I'm asking questions lol :-)

OpenStudy (anonymous):

another medal for you!

OpenStudy (callisto):

lol

OpenStudy (anonymous):

I like these molarity questions, but it's been a long time since I've had to do them.

OpenStudy (anonymous):

Hey, side question: If you have 0.1 Molar solution of NaOH and you want to get it to .05 M concentration, how do you figure out the volume of H20 you need to add?

OpenStudy (anonymous):

wait, I must have just answered my own question. Same method I guess.

OpenStudy (callisto):

sorry, was having lunch, yea, it's the same method..